OpenStudy (anonymous):

Clair bought three bars of soap and five sponges for $2.31. Steve bought five bars of soap and three sponges for $3.05. find the cost of each item

OpenStudy (anonymous):

3s+5p=2.31 5s+3p= 3.05 that's the equation

OpenStudy (anonymous):

What do you do to get the answer?

OpenStudy (anonymous):

you multiply numbers to each side to get one of the same number

OpenStudy (anonymous):

so u can do this: 5(3s+5p)= 5(2.31) 3(5s+3p)= 3(3.05)

OpenStudy (anonymous):

so its 15s+25p= 11.55 -15s-9p= -9.05 sorry i forgot its supposed to be -3(5s+3p)= -3(3.05) on the top

OpenStudy (anonymous):

then u cancel out both of the 15s so p=2.5 p is the variable for the sponges so one sponge is $2.50

OpenStudy (anonymous):

and then you just use that to get the price of the soap